OpenStudy (anonymous):

The Thomas family had an internet system for their family room installed. The installation fee was $90 and the monthly internet charge is 30$. If the Thomas family has paid $450 so far, for how many months have they had the internet? Show all work.

OpenStudy (anonymous):

\[90+30x=450\] \[30x=450-90=360\] \[x=\frac{360}{90}=4\]

OpenStudy (anonymous):

typo!

OpenStudy (anonymous):

\[x=\frac{360}{30}=12\]